# SequencePAR: Understanding Pedestrian Attributes via A Sequence Generation Paradigm

### v1

@@ Abstract
Current pedestrian attribute recognition (PAR) algorithms are developed based on multi-label or multi-task learning frameworks, which aim to discriminate the attributes using specific classification heads. However, these discriminative models are easily influenced by imbalanced data or noisy samples. Inspired by the success of generative models, we rethink the pedestrian attribute recognition scheme and believe the generative models may perform better on modeling dependencies and complexity between human attributes. In this paper, we propose a novel sequence generation paradigm for pedestrian attribute recognition, termed SequencePAR. It extracts the pedestrian features using a pre-trained CLIP model and embeds the attribute set into query tokens under the guidance of text prompts. Then, a Transformer decoder is proposed to generate the human attributes by incorporating the visual features and attribute query tokens. The masked multi-head attention layer is introduced into the decoder module to prevent the model from remembering the next attribute while making attribute predictions during training. Extensive experiments on multiple widely used pedestrian attribute recognition datasets fully validated the effectiveness of our proposed SequencePAR. The source code and pre-trained models will be released at https://github.com/Event-AHU/OpenPAR.
Index Terms:
@@
I Introduction
Pedestrian Attribute Recognition (PAR) [1, 2] involves utilizing a predefined set of attributes to estimate or infer semantic characteristics for pedestrian images, such as age, gender, dress, movement, etc. It plays an important role in many practical scenarios like intelligent video monitoring, autonomous driving, and pedestrian analysis. Pedestrian attribute recognition also plays a significant role in other pedestrian-related tasks, such as assisting in matching IDs for pedestrian re-identification tasks [3, 4, 5], object detection [6], natural language description-based pedestrian search or retrieval tasks [7, 8, 9], and analyzing and mining pedestrian attributes to predict pedestrian behavior intentions and possible action trajectories [10].
@@ I Introduction
Due to its irreplaceable important role, many deep learning based pedestrian attribute recognition algorithms have been proposed recently and made great progress. Specifically, numerous researchers have utilized Convolutional Neural Networks (CNNs) to acquire feature representations from pedestrian images. Subsequently, they employ fully connected layers [11, 12] to regress the responses of each attribute. To model the relations between human attributes, the Recurrent Neural Network (RNN) and the Graph Neural Network (GNN) are introduced to further enhance the CNN features [13, 14, 15]. Recently, the Transformer [16, 17] is also adapted for pedestrian attribute recognition [18, 19, 20, 21, 22] due to its strong ability on modeling the long-range global relations. For example, Cheng et al. propose the VTB [18] which achieves high-performance attribute recognition via multimodal Transformer based image-text fusion. The introduction of these algorithms has greatly promoted the development of the pedestrian attribute recognition field.
However, we find that these pedestrian attribute recognition algorithms basically adopt the framework of multi-label classification or multi-task learning, which belongs to the category of discriminative models. It is well known that these discriminative models are susceptible to the following factors:
  • Existing works directly utilize the CNNs/Transformers to predict attribute response scores which can be easily disturbed by the issue of data imbalance. A large number of negative samples may also lead to sparse attribute prediction due to the issue of data sparsity.
  • The annotation of human attributes is a challenging task (for example, some datasets [23] provide uncertain attribute annotations), and the annotated labels for the training data may contain noise or errors, which can affect the performance and reliability of the model.
  • The semantic connection between attributes is weak due to the simultaneous regression of each attribute.
As we all know, generative models can better model dependencies and complexity between data over discriminative models. It describes the relationships between multiple tasks through joint probability distributions and thus performs well in capturing the complexity and diversity of human attributes. The generative models can improve model performance and generalization by incorporating external knowledge or prior information. Therefore, a highly intuitive question worthy of consideration is: “Can we break away from the category of discriminative models and re-examine pedestrian attribute recognition from the perspective of generative models?"
@@
Inspired by the aforementioned observations and thinking, in this work, we propose a novel sequence generation paradigm for pedestrian attribute recognition, termed SequencePAR. A comparison between existing Transformer based pedestrian attribute recognition algorithms and our SequencePAR is illustrated in Fig. 1. The key insight of this work is that we formulate attribute recognition as an image captioning task, which can model the relations between human attributes by generating descriptive captions for human images. This generative model allows us to encapsulate the underlying connections between attributes. Furthermore, the burgeoning evidence from various large-scale models strongly indicates the exceptional performance of big language models in precisely this domain. In our practical implementation, given the pedestrian image, we first partition it into non-overlapping tokens and adopt the CLIP visual encoder [24] to get the visual tokens due to its strong feature expression ability and generalization. We also take the attribute descriptions as the input to capture the high-level semantic information. We adopt the word embedding to transform the attributes into text embeddings and concatenate them with text prompts as the attribute query token. Then, we feed the attribute query tokens into the masked multi-head attention layer, and the outputs are fed into the multi-head attention layer together with visual tokens for attribute generation. Note that, the normalization layers and feed-forward layers are also incorporated in the sequence generation decoder network. In the testing phase, the greedy search mechanism is adopted for attribute generation. An overview of our proposed SequencePAR framework can be found in Fig. 2.
@@
    Extensive experiments on multiple popular pedestrian attribute recognition datasets fully validated the effectiveness of our proposed SequencePAR.
@@
The rest of this paper is organized as follows: We review related works most related to our paper in Section II. The approach we proposed is mainly introduced in Section III, including the problem formulation, overview, input representation, sequence generation module, and loss function. In Section IV, we conduct extensive experiments and give an in-depth analysis of multiple benchmark datasets. We conclude this paper and propose possible research directions in Section V.
II Related Work
@@
In this section, we will give a brief introduction to the related works on Pedestrian Attribute Recognition, Transformer Networks and Sequence Generation. More related works can be found in the following survey [1] 111github.com/wangxiao5791509/Pedestrian-Attribute-Recognition-Paper-List.
II-A Pedestrian Attribute Recognition
Pedestrian attribute recognition methods can be categorized into various approaches, such as CNN, RNN, graphic model-based algorithms, attention mechanisms, and Transformer-based methods. Early CNN-based approaches have demonstrated promising results in pedestrian attribute recognition. For instance, MTCNN [11] proposed a multi-task learning approach, utilizing multiple CNNs to learn attribute-specific features while sharing knowledge among them. PGDM [25] predicts key points of the human body to leverage deformable body structure information for pedestrian attribute prediction, which captures pose variations and enhances attribute prediction performance. The RNN model is introduced into the PAR community to model the semantic association between human attributes. To be specific, Wang et al. [26] employed LSTM to establish high-level semantic dependencies between labels in pedestrian attribute recognition. By incorporating the previously predicted labels, visual features can be dynamically adjusted for the subsequent moments. Zhao et al. [27] employ convolutional LSTM to explore correlations between attribute groups. They also use recursive attention to leverage spatial locality within each group and attentional correlations between different attribute groups. Graphic models can be employed to establish connections or relationships between attributes, as well as between attributes and pedestrian components. Specifically, Chen et al. [28] integrate multiple features by combining them and feeding them into a classifier to obtain confidence levels. Subsequently, they employed conditional random fields (CRF) to explore and extract associations between attributes. The attention mechanism is widely used in deep neural networks and attribute recognition tasks. For example, Liu et al. introduce HydraPlus-Net [29], an architecture that leverages the inception-v2 [30] framework to generate multi-level feature maps. This approach enables the network to capture diverse pedestrian details and extract more comprehensive features.
Inspired by the success of Transformer networks in natural language processing and computer vision tasks, some researchers also adapted Transformers [16, 17, 31] for pedestrian attribute recognition. For instance, Weng et al. [32] utilize the Attribute Location Module(ALM) to obtain attribute-specific feature representations, and leverage the Transformer to enhance the connection between attribute-related features and the modeled attributes. Cheng et al. [18] approached pedestrian attribute recognition as a visual language task by converting the attribute list into text features. They fed both the visual features extracted from pedestrian images and the text features into a Transformer model, facilitating effective interaction between the modalities. Different from these works, we employ a generative framework that formulates pedestrian attribute recognition as a sequence generation problem. Through iterative attribute generation, our approach explicitly models the semantic relationships between attributes. This addresses the limitations of traditional pedestrian attribute recognition methods, which often overlook the modeling of relationships within attributes or rely on manual assumptions.
II-B Transformer Network
@@
Based on self-attention mechanisms, the Transformer [16] network first caused a significant stir in the field of natural language processing, demonstrating performance far superior to RNN/LSTM [33] series algorithms across various text-related tasks. Then, the pre-training techniques developed based on the Transformer further demonstrate its effectiveness, for example, the BERT [31], GPT series [34, 35], LLaMA [36], LaMDA [37], Baichuan-2 [38]. The remarkable success of the Transformer in natural language processing (NLP) has inspired researchers to explore its application in computer vision. Significant efforts have been made to introduce the Transformer into multiple fields. For example, the ViT [17] proposed by Dosovitskiy et al. is the first milestone of the Transformer-based foundation model. It splits the input image into token representations and introduces a class token before feeding it into the Transformer layers for classification. M. Cornia et al. [39] are pioneers in applying the Transformer model to the image captioning task. S. Herdade et al. [40] integrate spatial information into image captioning by fusing geometric attention weights into the attention mechanism.
@@
After that, the Transformer network is also introduced into the multi-modal community, especially for the vision-language-based pre-training and downstream tasks [41]. Many representative vision-language models are proposed one after another, such as the CLIP [24], ALIGN [42], LXMERT [43], ViLBERT [44]. Inspired by these works, we adopt the Transformer-based pre-trained models to extract the visual and textual representations. More importantly, we formulate the pedestrian attribute recognition task as a sequential text generation problem to better capture the semantic relations between various attributes.
II-C Sequence Generation Models
@@
Many high-level tasks involve generating sequential outputs, such as machine translation, image captioning, and speech recognition. These models usually follow the encoder-decoder framework and achieve this target via recurrent neural networks (RNN) or Transformer decoders. More in detail, the LSTM (Long Short-Term Memory) [33] is widely used for temporal information processing. Chen et al. [45] investigate the impact of different label orders when utilizing LSTMs for multi-label recognition tasks. Attention mechanisms are employed to address the issue of encoder-decoder models when dealing with fixed-length vectors. Lu et al. [46] introduce the concept of a visual sentinel which enables the model to dynamically determine whether to prioritize the visual image or the linguistic rule during different stages of the decoding process. Chen et al. [47] formulate the visual tracking as a sequence generation problem, which predicts object bounding boxes in an auto-regressive fashion. Chen et al. [48] cast object detection as a language modeling task conditioned on the observed pixel inputs and propose a simple and generic framework for object detection termed Pix2Seq. Li et al. [49] propose the BLIP-2 which bootstraps vision-language pre-training using a lightweight Querying Transformer. This work demonstrates that image-to-text generation can follow the introductions of natural language. In this work, we formulate pedestrian attribute recognition as a sequence generation problem instead of the standard multi-label classification problem.
III Methodology
@@
Figure 2: An overview of our proposed SequencePAR framework. Given the pedestrian image, we first partition it into non-overlapping patches and extract the visual tokens using pre-trained big models. Unlike previous works that directly regress the attribute scores utilizing a set of dense layers, in this work, we design a new generative paradigm that describes attributes sequentially. It is mainly achieved by a Transformer decoder, which takes the attribute query tokens and pedestrian image features as the input. More in detail, the masked multi-head attention layers are proposed to process the attribute query tokens. Note that, the masking mechanism ensures the model relies on contextual understanding for attribute prediction, preventing opportunistic memorization of subsequent attributes while predicting the current one. We adopt the multi-head cross-attention to fuse the visual and processed query tokens for attribute sequence generation.
III-A Problem Formulation
Pedestrian attribute recognition targets predicting human attributes $\mathcal{A}_{i}$, $i=\{1,2,...,N\}$, from a pre-defined attribute set $\mathcal{A}_{set}=\{\mathcal{A}_{1},\mathcal{A}_{2},...,\mathcal{A}_{M}\}$, $N\leq M$, based on a given pedestrian image $\mathcal{I}$. Existing deep learning-based attribute recognition models usually treat this task as a multi-label classification or multi-task learning problem [18, 15]. They usually adopt a backbone network to extract the deep features of the human image and learn to recognize the attributes using multi-category classifiers (MCC), i.e.,
@@ III-A Problem Formulation
where $VB$ denotes the visual backbone. Some researchers have proposed to recognize human attributes by combining pedestrian image features and attribute text representations.
@@
III-B Overview
In this work, we formulate the pedestrian attribute recognition task as a sequence generation problem similar to image captioning and machine translation. Given the pedestrian image, we first partition and project it into token representations as the input of the Transformer network. To capture the high-level semantic information of attributes, we also take the raw attributes as the input and utilize the embedded attribute features as the attribute query tokens. The attribute query tokens are processed by masked self-attention layers and further fused with pedestrian image features via cross-attention layers. Such a joint processing of image and text features facilitates the effective integration of multi-modal data for attribute prediction. Finally, the resulting features pass through a linear layer followed by a Softmax function to generate predicted probabilities for each attribute. Therefore, we can get the final attributes via rearrangement. An overview of our proposed SequencePAR framework can be found in Fig. 2. In the following subsections, we will focus on each detailed procedure to help the readers better understand our proposed SequencePAR model.
III-C Network Architecture
Image Encoding Network.  Given a pedestrian attribute recognition dataset $\mathcal{D}=\{(X_{i},A_{i})|i=\{0,1,2,...,N\}\}$, where $X_{i}$ and $A_{i}=\{a_{i}^{1},a_{i}^{2},...,a_{i}^{M}\}$ represents pedestrian images and corresponding attribute labels, respectively. $N$ is the number of pedestrian images in the dataset and $M$ represents the number of labeled attributes. We first split the image into $256$ non-overlapping patches $P_{i}\in\mathbb{R}^{14\times 14}$, then, project them into token representation $T_{i}\in\mathbb{R}^{256\times 1024}$ using one convolutional layer (kernel size 14 $\times$ 14). Following the standard vision Transformer networks, we also introduce the position encoding $PE\in\mathbb{R}^{257\times 1024}$ to capture the spatial information of split patches. The token features and position features are added together as the input of the vision Transformer network. In this paper, we adopt the pre-trained vision encoder of multi-modal big model CLIP [24] (ViT-L/14 is adopted) to extract the feature representations $\hat{T}_{I}\in\mathbb{R}^{257\times 768}$. The core components of ViT-L/14 are the Transformer layers and each layer consists of multi-head self-attention layers, normalization layers, feed-forward layers, and skip connections. Among them, the detailed procedure of multi-head self-attention is the self-attention operation, which can be written as:

### v2

@@ Abstract
Current pedestrian attribute recognition (PAR) algorithms use multi-label or multi-task learning frameworks with specific classification heads. These models often struggle with imbalanced data and noisy samples. Inspired by the success of generative models, we propose Sequence Pedestrian Attribute Recognition (SequencePAR), a novel sequence generation paradigm for PAR. SequencePAR extracts pedestrian features using a language-image pre-trained model and embeds the attribute set into query tokens guided by text prompts. A Transformer decoder generates human attributes by integrating visual features and attribute query tokens. The masked multi-head attention layer in the decoder prevents the model from predicting the next attribute during training. The extensive experiments on multiple PAR datasets validate the effectiveness of SequencePAR. Specifically, we achieve 84.92%, 90.44%, 90.73%, and 90.46% in accuracy, precision, recall, and F1-score on the PETA dataset.
keywords:
@@
journal: Pattern Recognition
1 Introduction
Pedestrian Attribute Recognition (PAR) [30] involves utilizing a predefined set of attributes to estimate or infer semantic characteristics for pedestrian images, such as age, gender, dress, movement, etc. It plays an important role in many practical scenarios like intelligent video monitoring, autonomous driving, and pedestrian analysis. Pedestrian attribute recognition also plays a significant role in other pedestrian-related tasks, such as assisting in matching IDs for pedestrian re-identification tasks [40], object detection [38], natural language description-based pedestrian search or retrieval tasks [23], and analyzing and mining pedestrian attributes to predict pedestrian behavior intentions and possible action trajectories [18].
@@ 1 Introduction
Due to its irreplaceable important role, many deep learning-based pedestrian attribute recognition algorithms have been proposed recently and made great progress. Specifically, numerous researchers have employed Convolutional Neural Networks (CNNs) to extract feature representations from pedestrian images, followed by the use of fully connected layers to regress the attribute responses [26]. Additionally, models like Recurrent Neural Networks (RNNs) and Graph Neural Networks (GNNs) have been introduced to capture the complex relationships between human attributes, enhancing the CNN features [7]. More recently, Transformers [29], known for their ability to model long-range dependencies, have been adapted for pedestrian attribute recognition tasks [2]. For example, Cheng et al. propose VTB [2], which achieves high performance via multimodal Transformer-based image-text fusion. These advances have undoubtedly propelled the field of pedestrian attribute recognition.
However, despite these advancements, existing pedestrian attribute recognition algorithms typically operate within the multi-label classification framework or multi-task learning, which are categorized as discriminative models. It is well known that discriminative models are susceptible to various challenges, including data imbalance and sparsity, noisy annotations, and insufficient semantic relationship mining, which may constrain their performance.
To address the limitations of discriminative models, we propose a shift in perspective: “Can we break away from the category of discriminative models and re-examine pedestrian attribute recognition from the perspective of generative models?” Generative models, by their nature, can better capture complex relationships and dependencies between attributes. Unlike discriminative models, which focus on classifying or regressing individual attributes independently, generative models transform $N$ independent binary classification problems into a sequence joint probability model $P\left(A_{1},A_{2},\ldots,A_{N}\right)=\prod{P\left(A_{i}\middle|A<i\right)}$, dynamically constructing conditional probability relationships between attributes during the sequence generation process. This explicit modeling approach enables tail attributes to infer from the contextual information of head attributes (e.g., the ”skirt” attribute provides strong semantic constraints for predicting ”female”), effectively alleviating the issue of poor learning of tail features due to isolated classification in traditional methods. Regarding noise robustness, during the autoregressive generation process, the prediction of each attribute is constrained by the semantic constraints of previously generated attributes. For example, when a noisy sample incorrectly labels ”short-sleeve + skirt” as male, the model can automatically correct the subsequent gender prediction through the already generated ”skirt” attribute (strongly associated with female). This reasoning mechanism constructs a dynamic error tolerance space. These characteristics make generative models particularly suited for pedestrian attribute recognition, where relationships between attributes play a crucial role in achieving high performance.
@@
Inspired by the aforementioned observations and thinking, in this work, we propose a novel sequence generation paradigm for pedestrian attribute recognition, termed SequencePAR. A comparison between existing Transformer-based pedestrian attribute recognition algorithms and our SequencePAR is illustrated in Figure 1. The key insight of this work is that we formulate attribute recognition as an image captioning task, which can model the relations between human attributes by generating descriptive captions for human images. This generative model allows us to encapsulate the underlying connections between attributes. Furthermore, the burgeoning evidence from various large-scale models strongly indicates the exceptional performance of big language models in precisely this domain. In our practical implementation, given the pedestrian image, we first partition it into non-overlapping tokens and adopt the CLIP visual encoder [24] to get the visual tokens due to its strong feature expression ability and generalization. We also take the attribute descriptions as the input to capture the high-level semantic information. We adopt the word embedding to transform the attributes into text embeddings and concatenate them with text prompts as the attribute query token. Then, we feed the attribute query tokens into the masked multi-head attention layer, and the outputs are fed into the multi-head attention layer together with visual tokens for attribute generation. Note that, the normalization layers and feed-forward layers are also incorporated in the sequence generation decoder network. In the testing phase, the greedy search mechanism is adopted for attribute generation. An overview of our proposed SequencePAR framework can be found in Figure 2.
@@
    Extensive experiments on multiple popular pedestrian attribute recognition datasets fully validated the effectiveness of our proposed SequencePAR. The source code and pre-trained models will be released at https://github.com/Event-AHU/OpenPAR.
@@
The rest of this paper is organized as follows: We review related works most related to our paper in Section 2. The approach we proposed is mainly introduced in Section 3, including the problem formulation, overview, input representation, sequence generation module, and loss function. In Section 4, we conduct extensive experiments and give an in-depth analysis of multiple benchmark datasets. We conclude this paper and propose possible research directions in Section 5.
2 Related Works
@@
In this section, we will give a brief introduction to the related works on Pedestrian Attribute Recognition, Transformer Networks, and Sequence Generation. More related works can be found in the following survey [30] 111github.com/wangxiao5791509/Pedestrian-Attribute-Recognition-Paper-List.
2.1 Pedestrian Attribute Recognition
Pedestrian attribute recognition methods can be categorized into various approaches, such as CNN, RNN-based algorithms, attention mechanisms, and transformer-based methods. Early CNN-based approaches have demonstrated promising results in pedestrian attribute recognition. For instance,
DAFL [11] enables the model to learn attribute-independent visual features through a group-shared attention library, thereby enhancing the discriminative capability of the predictions. Shen et al. [26] proposed a method that leverages multi-scale priors and attribute-space priors to enable the model to accurately localize attribute regions. Zhou et al. [39] proposed extracting attribute-specific cues to mitigate the model’s over-reliance on attribute co-occurrence patterns in the dataset, thereby enabling the model to better adapt to imbalanced distributions. The RNN model is introduced into the PAR community to model the semantic association between human attributes. By incorporating the previously predicted labels, visual features can be dynamically adjusted for the subsequent moments. GRL [36] employs an RNN to capture the relationships and mutual exclusivities among the attributes. The attention mechanism is widely used in deep neural networks and attribute recognition tasks. For example, Liu et al. introduce HydraPlus-Net [19], an architecture that leverages the multi-level feature maps to capture diverse pedestrian details and extract more comprehensive features.
Inspired by the success of Transformer networks in natural language processing and computer vision tasks, some researchers also adapted Transformers [29, 6, 13] for pedestrian attribute recognition. For instance, Wu et al. [33] introduce perturbations into the attention maps during training to reduce the model’s focus on non-attribute regions, thereby enhancing its ability to capture attribute-specific areas. Cheng et al. [2] approached pedestrian attribute recognition as a visual language task by converting the attribute list into text features. They fed both the visual features and the text features into a Transformer model, facilitating effective interaction between the modalities.
However, discriminative models face various challenges that may constrain their performance:
  • 1.
    Data Imbalance and Sparsity: Models that directly use CNNs or Transformers to predict attribute responses often struggle with data imbalance, where a large number of negative samples can lead to sparse attribute predictions. This is particularly problematic in pedestrian attribute recognition, where certain attributes are underrepresented in the dataset.
  • 2.
    Noisy Annotations: Pedestrian attribute annotation is a complex and error-prone task. Some datasets [16] contain uncertain or noisy attribute labels, which can significantly impact model reliability and performance.
  • 3.
    Weak Semantic Relationships: In discriminative models, where attributes are regressed independently, the semantic relationships between different attributes are often weak or underexplored, leading to suboptimal modeling of attribute dependencies.
To address these issues, we employ a generative framework that formulates pedestrian attribute recognition as a sequence generation problem. Through iterative attribute generation, our approach explicitly models the semantic relationships between attributes. This addresses the limitations of traditional pedestrian attribute recognition methods, which often overlook the modeling of relationships within attributes or rely on manual assumptions.
2.2 Transformer Network
@@
Based on self-attention mechanisms, the Transformer [29] network first caused a significant stir in the field of natural language processing, demonstrating performance far superior to RNN//LSTM series algorithms across various text-related tasks. Then, the pre-training techniques developed based on the Transformer further demonstrate its effectiveness, for example, the BERT [13]. The remarkable success of the Transformer in natural language processing (NLP) has inspired researchers to explore its application in computer vision. Significant efforts have been made to introduce the Transformer into multiple fields. For example, the ViT [6] proposed by Dosovitskiy et al. is the first milestone of the Transformer-based foundation model. It splits the input image into token representations and introduces a class token before feeding it into the Transformer layers for classification. M. Cornia et al. [3] are pioneers in applying the Transformer model to the image captioning task.
@@
After that, the Transformer network is also introduced into the multi-modal community, especially for the vision-language-based pre-training and downstream tasks. Many representative vision-language models are proposed one after another, such as the CLIP [24]. Inspired by these works, we adopt the Transformer-based pre-trained models to extract the visual and textual representations. More importantly, we formulate the pedestrian attribute recognition task as a sequential text generation problem to better capture the semantic relations between various attributes.
2.3 Sequence Generation Models
@@
Many high-level tasks involve generating sequential outputs, such as machine translation, image captioning, and speech recognition. These models usually follow the encoder-decoder framework and achieve this target via recurrent neural networks (RNN) or Transformer decoders. More in detail, the LSTM is widely used for temporal information processing. Chen et al. [1] formulate the visual tracking as a sequence generation problem, which predicts object bounding boxes in an auto-regressive fashion. Li et al. [17] propose the BLIP-2 which bootstraps vision-language pre-training using a lightweight Querying Transformer. This work demonstrates that image-to-text generation can follow the introductions of natural language. Peng et al. [22] propose an unbiased VQA framework that highlights the potential of generative formulations in visual question answering, further demonstrating the effectiveness of sequence generation paradigms in multimodal reasoning tasks. In this work, we formulate pedestrian attribute recognition as a sequence generation problem instead of the standard multi-label classification problem.
3 Methodology
@@
3.1 Problem Formulation
Pedestrian attribute recognition targets predicting human attributes $\mathcal{A}_{i}$, where $i=\{1,2,...,N\}$, from a pre-defined attribute set $\mathcal{A}_{set}=\{\mathcal{A}_{1},\mathcal{A}_{2},...,\mathcal{A}_{M}\}$, $N\leq M$, based on a given pedestrian image $\mathcal{I}$. Existing deep learning-based attribute recognition models usually treat this task as a multi-label classification or multi-task learning problem [2, 8]. They usually adopt a backbone network to extract the deep features of the human image and learn to recognize the attributes using multi-category classifiers (MCC), i.e.,
@@ 3.1 Problem Formulation
where $VB$ denotes the visual backbone. Some researchers have proposed recognizing human attributes by combining pedestrian image features and attribute text representations.
@@
Figure 2: An overview of our proposed SequencePAR framework. Given a pedestrian image, we first partition it into non-overlapping patches and extract visual tokens using pre-trained large models. Unlike previous approaches that directly regress attribute scores through dense layers, our method introduces a novel generative paradigm that sequentially generates attribute descriptions.
3.2 Overview
In this work, we formulate the pedestrian attribute recognition task as a sequence generation problem similar to image captioning and machine translation. Given the pedestrian image, we first partition and project it into token representations as the input of the Transformer network. To capture the high-level semantic information of attributes, we also take the raw attributes as the input and utilize the embedded attribute features as the attribute query tokens. The attribute query tokens are processed by masked self-attention layers and further fused with pedestrian image features via cross-attention layers. Such a joint processing of image and text features facilitates the effective integration of multi-modal data for attribute prediction. Finally, the resulting features pass through a linear layer followed by a Softmax function to generate predicted probabilities for each attribute. Therefore, we can get the final attributes via rearrangement. An overview of our proposed SequencePAR framework can be found in Figure 2. In the following subsections, we will focus on each detailed procedure to help the readers better understand our proposed SequencePAR model.
3.3 Network Architecture
Image Encoding Network.  Given a pedestrian attribute recognition dataset $\mathcal{D}=\{(X_{i},A_{i})|i=\{0,1,2,...,N\}\}$, where $X_{i}$ and $A_{i}=\{a_{i}^{1},a_{i}^{2},...,a_{i}^{M}\}$ represents pedestrian images and corresponding attribute labels, respectively. $N$ is the number of pedestrian images in the dataset and $M$ represents the number of labeled attributes. We first split the image into $256$ non-overlapping patches $P_{i}\in\mathbb{R}^{14\times 14}$, then, project them into token representation $T_{i}\in\mathbb{R}^{256\times 1024}$ using one convolutional layer (kernel size 14 $\times$ 14). Following the standard vision Transformer networks, we also introduce the position encoding $PE\in\mathbb{R}^{257\times 1024}$ to capture the spatial information of split patches. The token features and position features are added together as the input of the vision Transformer network. In this paper, we adopt the pre-trained vision encoder of multi-modal big model CLIP [24] (ViT-L//14 is adopted) to extract the feature representations $\hat{T}_{I}\in\mathbb{R}^{257\times 768}$. The core components of ViT-L//14 are the Transformer layers and each layer consists of multi-head self-attention layers, normalization layers, feed-forward layers, and skip connections. Among them, the detailed procedure of multi-head self-attention is the self-attention operation, which can be written as: